A consumer buys 80 units of a good at a price of rupees 5 per unit. Suppose price elasticity of demand is -2 at what price will he buy 64 units?


1
Expert's answer
2020-12-08T07:50:38-0500

"\bold {Answer}"


"Price = 4.50 \space rupees \space per \space unit"


"\bold {Solution}"

"PED = \dfrac {\%∆ \space Quantity}{\%∆ \space Price}"


"\% ∆ \space Quantity = \dfrac {80-64}{80}×100\%"

"=\dfrac {16}{80}×100\%"


"\bold {= 20\%}"


Let "x" be the price when "Q= 64 \space units"


"\%∆\space Price = \dfrac {x -5}{5}×100\%"


"\bold {= 20(x-5)\%}"


Therefore, "-2 = \dfrac {20\%}{20(x-5)\%}"


"=> -2(x-5) = 1"

"=> -2x + 10 = 1"

"=> -2x = 1-10"

"-2 x = -9"

Dividing both sides by "-2" gives:

"\bold {x = 4.50}"


Therefore, when the consumer buy 64 units, the price is 4.50 rupees per unit.